OpenStudy (anonymous):

Prove by contradiction that there is no rational number r such that r^3 + r + 1 = 0. Assume r^3 + r + 1 = 0. Let r = a/b. Then (a/b)^3 + a/b + 1/1 = 0 a^3/b^3 + a/b + 1/1 = 0 (ab^2 + a^3 + b^3)/b^3 = 0 ab^2 + a^3 + b^3 = 0 I'm not seeing the proof... Help?

OpenStudy (perl):

better to ask, does that equation have a rational x intercept?

OpenStudy (perl):

use the rational roots theorem on f(r) = r^3 + r + 1

ganeshie8 (ganeshie8):

you know the proof by contradiction for proving sqrt(2) is irrational right ?

OpenStudy (perl):

if f (p/q) = 0 , then p divides 1 and q divides 1. therefore the potential candidates for p/q can only be 1, -1 f(1) = 3 f(-1) = -1 , so both of these are not rational roots. youre done

OpenStudy (perl):

i think that works, ill have to think about it

ganeshie8 (ganeshie8):

thats not proof by contradiction perl even though it works

OpenStudy (perl):

there might be a way to present it as a proof by contradiction. if you assume there exists p/q such that f(p/q) = 0, for f(r) = r^3 + r + 1

OpenStudy (perl):

then the rational roots theorem says p/q = {1, -1} . but both of these f(1) , f(-1) does not equal to zero. contradiction

OpenStudy (perl):

its a contradiction because we assumed there does exist a solution f(p/q) = 0

OpenStudy (ikram002p):

is zero a rational number ?

OpenStudy (perl):

0 = 0/1 , so 0 is rational

ganeshie8 (ganeshie8):

my point is rational root theorem should not be used in a proof by contradiction

OpenStudy (perl):

why not?

ganeshie8 (ganeshie8):

if u are using rational root theorem anyways, you could present it as direct proof itself... no need to twist

ganeshie8 (ganeshie8):

its not what is expected in a contradiction proof. it works though

OpenStudy (perl):

ok there might be a way to show it using number theoretic properties

ganeshie8 (ganeshie8):

looks hard without rational root thereom :O

OpenStudy (ikram002p):

ohk r^3+r+1=0 r^3+r=-1 r(r^2+1)=-1 so four cases case 1:- r=-1 and r^2+1=1...r^2=0 ignore case 2:- r=1 , r^2+1=1 r^2=2 ignore case 3:- r=a/b , and r^2+1=-a/b case 3:- r=-a/b and r^2+1=a/b lets work on case 3 and 4

OpenStudy (ikram002p):

sorry i made a typo xD case 3:- r=a/b , and r^2+1=-b/a case 4:- r=-a/b and r^2+1=b/a we might see also other two cases case 5:- r=b/a , and r^2+1=-a/b case 4:- r=-b/a and r^2+1=a/b

ganeshie8 (ganeshie8):

case 3:- r=a/b , and r^2+1=-b/a right ?

OpenStudy (ikram002p):

yep :)

ganeshie8 (ganeshie8):

doesn't case3 cover all other cases ?

ganeshie8 (ganeshie8):

just let a, b to be integers

OpenStudy (ikram002p):

ok thats work i just create this up lol im gonna see where it gonna end :) now we assume a/b is rational why ? cuz that what we need to proof , if its not rational like pi/something then we are done :)

OpenStudy (ikram002p):

now case 3 :- (PS :a and b are integers s.t b>=a) r=a/b r^2+1=-b/a r^2=-b/a-1 =-(b+a)/a (a/b)^2=-(b+a)/a which is a contradiction

OpenStudy (ikram002p):

same with other cases xD

ganeshie8 (ganeshie8):

why is the last line a contradiction ?

OpenStudy (ikram002p):

last line u have one side is positive and the other is negative

ganeshie8 (ganeshie8):

Oh you're assuming a,b are positive to start with

OpenStudy (ikram002p):

yeah thats why i gave all cases , sorry for not mention that xD

ganeshie8 (ganeshie8):

i see, so we are done if we show all cases lead to contradiction

OpenStudy (ikram002p):

yep , i still need to write the rest

OpenStudy (ikram002p):

ok case 4 r=-a/b and r^2+1=b/a r^2=b/a-1 blah blah a^3-b^3=-ab^2 which is a contradiction since the only integer solution are 0,0

ganeshie8 (ganeshie8):

got it :)

OpenStudy (ikram002p):

oh , got other method ? :O

ganeshie8 (ganeshie8):

nope im saying i got ur method :)

OpenStudy (ikram002p):

ohkk ...

OpenStudy (ikram002p):

ok so these are only cases other cases are like this b>a r=b/a r^2+1=-a/b contradiction r^2+1>0

OpenStudy (ikram002p):

case 6 b>a r=-b/a r^2=(b/a)^2>r r^2+1=(b/a)^2+1>r+1 r^2+1=a/b contradiction

OpenStudy (ikram002p):

done :O

myininaya (myininaya):

This thing finally loaded for me. I sent @ganeshie8 a message thinking I would never be able to get back here. I will post it.

myininaya (myininaya):

This is what it said: Hey I can't get back to that question. Submit this if you think it is a good idea. I will only put in some things since I'm limited here. r^3+r+1+r^2-r^2=0 where r=a/b is a rational. Let f(r)=r^3+r^2+r+1. Assuming r=a/b we have r^3+r^2+r+1=(a+b)(a^2+b^2)=0. So a=-b. So r=-1. So r+1 is a factor of f(r). So we have r^3+r+1+r^2-r^2 can be written as (r+1)(r^2+1)-r^2=0. This can be written as r=-1/(r^2+1) which is not true for any rational r. Therefore r^3+r+1 has no rational zeros.

OpenStudy (ikram002p):

nice :)

ganeshie8 (ganeshie8):

wow! looks very neat r=-1/(r^2+1) is not possible is this because , left side and right side can never have same signs ?

ganeshie8 (ganeshie8):

oh no, when r is negative my reason fails.. so i think u have a better reason

OpenStudy (ikram002p):

r=-1/(r^2+1) r(r^2+1)=-1 i think this is what we nee to prove lol xD

OpenStudy (ikram002p):

ok i guess this leads to a/b=-1/((a^2+b^2)/b^2)=-b^2/a^2+b^2 a/b=-b^2/a^2+b^2 a(a^2+b^2)=-b^3 a^3+b^3+ab^2=0 has no integer solution exept 0,0

ganeshie8 (ganeshie8):

i have somehting like this in head: r^3+r+1+r^2-r^2=0 where r=a/b is a rational. (a+b)(a^2+b^2)=a^2b if a, b are odd then left side is even, right side is odd - contradiction if a, b are of opposite parity, left side is odd and right side is even - contradiction \(\square \)

OpenStudy (ikram002p):

nice also hmm

OpenStudy (ikram002p):

but but second case hmm

OpenStudy (ikram002p):

ok it works :S

ganeshie8 (ganeshie8):

how about this : a^3 + ab^2 + b^3 = 0 a(a^2+b^2) = -b^3 since a, b are coprime, a|1 similarly we can show that b|1

OpenStudy (ikram002p):

proof art lol

ganeshie8 (ganeshie8):

thats just a copy of rational root theorem proof ;p http://en.wikipedia.org/wiki/Rational_root_theorem

OpenStudy (ikram002p):

haha , i liked rational theorem thingy :O

OpenStudy (ikram002p):

i dont think that i would write such a beautiful or neat proof like that xD

myininaya (myininaya):

and yeah I guess \[-r=\frac{1}{r^2+1} \text{ is only obvious if } r<0\] Also I think a lot of that work I did was unnecessary :p.

myininaya (myininaya):

\[r^3+r+1=0 \\ r(r^2+1)+1=0 \\ r(r^2+1)=-1 \\ r=\frac{-1}{r^2+1} \\ -r=\frac{1}{r^2+1}\]

myininaya (myininaya):

i mean the equality is only possible if we have r<0* unless I can think of way to show it is not possible